Given the Circuit below: V1 R1 200 Ω 200 V R2 500 Ω R3 400 Ω V2 R4 300 Ω 400 V Calculate the Currents (Ii, I2, I5, 14) through each resistor
Answer #1

V1 R1 200 Ω 200 V R2 500 Ω R3 400 Ω R4 V2 in 300 Ω 400 V

since R2 and R3 are in parallel

i2 R2 = i3 R3

i2 (500) = i3 (400)

i3 = 1.25 i2

using KCL :

i1 + i4 = i2 + i3

i1 + i4 = 1.25 i2 + i2

i1 + i4 = 2.25 i2                                          eq-1

using KVL in loop ABCHA

i1 R1 - V1 + i2 R2 = 0

200 i1- 200 + 500 i2 = 0

i1= (200 - 500 i2 )/200                                            eq-2

using KVL in loop EDGFE

i3 R3 - V2 + i4 R4 = 0

400 i3- 400 + 300 i4 = 0

i4 = (400 - 400 i3 )/300                                                    eq-3

using eq-1 , eq-2 and eq-3

(200 - 500 i2 )/200 + (400 - 400 i3 )/300 = 2.25 i2  

(200 - 500 i2 )/200 + (400 - (400) (1.25) i2)/300 = 2.25 i2  

i2 = 0.364 A

i3 = 1.25 i2 = 1.25 x 0.364 = 0.455 A

i1= (200 - 500 i2 )/200 = (200 - 500 (0.364))/200 = 0.09 A

i4 = (400 - 400 i3 )/300   = (400 - 400 (0.455))/300 = 0.727 A